OpenStudy (anonymous):

f(x)=(x+6)^2 <= 0 the inequality has __________ solution(s) (type a whole number) OR the inequality has infinitely many solutions.

OpenStudy (skullpatrol):

Can the square of a number ever be < 0

OpenStudy (skullpatrol):

$$(a)^2<0$$

OpenStudy (anonymous):

no its \[f(x)=(x+6)^{2}\le0 \]

OpenStudy (skullpatrol):

Replace x+6 with a

OpenStudy (skullpatrol):

Can the square of a number ever be < 0

OpenStudy (anonymous):

it can but it can also be equal to 0

OpenStudy (skullpatrol):

Just think about one case at a time. Can $$a^2<0$$ yes or no?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What number squared is <0

OpenStudy (skullpatrol):

@flutterflies ?

OpenStudy (anonymous):

what do u mean what number squared? u mean (x+6)?

OpenStudy (skullpatrol):

I mean replace x+6 with a .. so $$(x+6)^2 = (a)^2 < 0$$

OpenStudy (anonymous):

so (x+6)

OpenStudy (skullpatrol):

What grade are you in?

OpenStudy (anonymous):

what does that have to do with the problem?????? i am just confuse on what u are trying to say. are u telling me to replace (x+6)^2 for a or for a^2

OpenStudy (skullpatrol):

I need to know what level of mathematical sophistication to talk to you with.

OpenStudy (anonymous):

i am taking an algebra class so just use a level of mathematical for algebra ;)

OpenStudy (skullpatrol):

What grade?

OpenStudy (anonymous):

level of mathematical sophistication should be around algebra..still not sure why you need the grade. you can just use the terms in algebra.

OpenStudy (skullpatrol):